A roof consists of steel beams/girders supporting 1-way slabs as shown. The weight of the slab, beams, and girders is estimatPlease show all calculations so I can follow along. The answer should be 76.1kN.
Answer #1

Tributary width of Beam 3=4/2=2 m

Load on beam 3=2*5=10 kN/m

Span of Beam 3=9 m

Load on C1 from beam 3=10*9/2=45 kN

Tributary widt of beam 1=1 m

Tributary width of beam 2=1+1=2 m

Load on Beam 1=1*5=5 kN/m

Load on beam2=2*5=10 kN/m

Load on girder 1 from beam 1=5*4/2=10 kN

Load on girder 1 from beam 2=10*4/2=20 kN

Load on girder 2 from girder 1=((10*5)+(20*7))/9=21.11 kN

Load on C1 from girder 2=21.1 kN

Total load on C1=45+21.1=66.1 kN